You are going to deposit $3,100 in an account that pays .57 percent interest per quarter. How much will you have in 7 years? Multiple Choice $3,655.50

$3,614.18

$3,634.78

$3,629.92

$3,638.82
Answer #1

PV =3100
Number of Quarters =7*4 =28
FV =PV*(1+r)^n =3100*(1+0.57%)^28 =3634.78 Option b is correct option